OpenStudy (angeleyes108):

CHECK MY ANSWER: The volume of a sphere is 4,000pi m3. What is the surface area of the sphere to the nearest square meter?

OpenStudy (angeleyes108):

Choices: A: 2614 B: 3836 C: 1307 D: 794

OpenStudy (angeleyes108):

My Answer: C

OpenStudy (anonymous):

SA (surface Area)=\[4\times3.14(pie)\times radius ^{2}\]

OpenStudy (angeleyes108):

All I have is the volume.

OpenStudy (angeleyes108):

i did it differently

OpenStudy (anonymous):

so then reverse it V= \[3/4\times 3.14(pie)\times radis ^{3}\]

OpenStudy (anonymous):

you get it?

OpenStudy (angeleyes108):

I did this: \[r^{3} = V \times \frac{ 3 }{ 4 }\pi = 3\times4000\pi \times \frac{ 3 }{ 4 }\pi = 3000cm^3 = 14.42 m\] and then plugged it in with 14.42 as radius

OpenStudy (angeleyes108):

am i right in doing that?

OpenStudy (anonymous):

i think

OpenStudy (angeleyes108):

ok thanks :)